You purchased 258 shares of stock in XYZ Company for $39 per share one year ago. You just sold your shares for $46 per share. You paid a 6% commission on each transaction. What is your rate of return for the year?

Express your % answer in decimal format: 4.5% = 0.045
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer
Statementshowing Computations
Paticulars Amount
Sale Value = 258*46*.94            11,155.92
Purchase cost = 258*39*1.06            10,665.72
Return                  490.20
rate of return = 490.20/10,665.72 4.60%